1.    A 1.000-g sample of ultrapure silicon was treated with an oxidizing agent and the resulting SiO2, after proper collection and drying, weighed 2.1387-g. Calculate the %Si in the sample.

 



1
Expert's answer
2022-04-26T09:05:05-0400

Solution:

analyte = ultrapure silicon (Si)

precipitate = SiO2




FW Si = 28.0855 g/mol

FW SiO2 = 60.08 g/mol

a = 1

b = 1

Therefore,

GF = (28.0855 / 60.08) × (1/1) = 0.46747


GF = Grams of analyte / Grams of precipitate

Therefore,

Grams of analyte = GF × Grams of precipitate

Grams of Si = GF × Grams of SiO2 = (0.46747) × (2.1387 g) = 0.9998 g


%Si = (Mass of Si / Mass of sample) × 100%

%Si = (0.9998 g / 1.000 g) × 100% = 99.98%

%Si = 99.98%


Answer: The %Si in the sample is 99.98%